sec A tan B+tan A secB = 2 then,
sec A sec B+tan A tan B = k.
then  k value

Grade:11

1 Answers

Aditya Gupta
3 years ago
k^2= sec^2a*sec^2b + tan^2a*tan^2b + 2seca*secb*tana*tanb
and 2^2= sec^2a*tan^2b + tan^2a*sec^2b + 2seca*secb*tana*tanb
subtract both eqns
k^2 – 4= sec^2a(sec^2b – tan^2b) + tan^2a(tan^2b – sec^2b)
now use identity tan^2x+1= sec^2x
k^2 – 4= sec^2a – tan^2a= 1
or k^2= 5
or k= ± sqrt(5)
KINDLY APPROVE :))